Alfred obtained a mark of 50% in his half yearly exam. He decided to go to private tuition with the hope of doing better in the annual exams. Alfred was rewarded because his mark went up to 70%. Find the percentage increase.
Answer by LinnW(1048)

70/50 - 1 = 140/100 - 1
1.40 - 1 = .40
So the increase is 40 percent